OpenStudy (anonymous):

Find the sum of a finite geometric sequence from n = 1 to n = 6, using the expression −2(5)n − 1

OpenStudy (anonymous):

@hwyl can you help?

OpenStudy (anonymous):

\[\sum_{n=1}^6-2(5)^{n-1}\]?

OpenStudy (anonymous):

i mean is that the question?

OpenStudy (anonymous):

i think so.. i don't remember how to do this

OpenStudy (anonymous):

pull the minus two out first, deal with it at the end

OpenStudy (anonymous):

\[\sum_{n=1}^6-2(5)^{n-1}=-2\sum_{n=1}^6(5)^{n-1}\]

OpenStudy (anonymous):

then \[\sum_{n=1}^6(5)^{n-1}=\frac{5^6-1}{5-1}\] compute that, multiply the restult by \(-2\)